Flex Co. just paid total dividends of $1,075,000 and reported additions to retained earnings of $3,225,000. The company has 715,000 shares of stock outstanding and a benchmark PE of 17.3 times. What stock price would you consider appropriate?
Answer #1

Current Year's Earnings = Dividends Paid + Addition to Retained Earnings

= 1075000 + 3225000 = $4300000

Earnings per Share = Earnings/Number of Shares

= 4300000/715000 = $6.014

Stock Price = Earnings per Share*PE Ratio

= 6.014*17.3 = $104.04